OpenStudy (anonymous):

Part A: Solve A = 9/2 (x + 23) for x. Part B: Determine the value of x when A = 108. Part C: Solve -np - 90 > 30 for n. Show your work.

OpenStudy (anonymous):

@ganeshie8 Help me on this one please??

ganeshie8 (ganeshie8):

\(\large A = \frac{9}{2} (x + 23) \)

ganeshie8 (ganeshie8):

to solve x, you need to isolate it you need to get x all by itself

ganeshie8 (ganeshie8):

wat shud be ur first step ? :)

OpenStudy (anonymous):

I can't do it with fractions :/ I have no clue

ganeshie8 (ganeshie8):

me too hate fractions, so lets get rid of them first multiply both sides wid 2

ganeshie8 (ganeshie8):

\(\large A = \frac{9}{2} (x + 23) \) multiply both sides wid 2 \(\large 2 \times A = 2 \times \frac{9}{2} (x + 23) \)

ganeshie8 (ganeshie8):

\(\large A = \frac{9}{2} (x + 23) \) multiply both sides wid 2 \(\large 2 \times A = 2 \times \frac{9}{2} (x + 23) \) \(\large 2A = 9(x + 23) \)

ganeshie8 (ganeshie8):

fractions are gone !

OpenStudy (anonymous):

Oh thank goodness xD So then what do i do?

ganeshie8 (ganeshie8):

next divide 9 both sides

ganeshie8 (ganeshie8):

\(\large A = \frac{9}{2} (x + 23) \) multiply both sides wid 2 \(\large 2 \times A = 2 \times \frac{9}{2} (x + 23) \) \(\large 2A = 9(x + 23) \) divide 9 both sides \(\large \frac{2A}{9} = x + 23\)

ganeshie8 (ganeshie8):

see, \(x\) is looking bit lean now it has oly 23 attached to it now get rid of it aswell

ganeshie8 (ganeshie8):

wat must we do... to rip off 23 from x

OpenStudy (anonymous):

Umm... and then am I supposed to subtract 23 from both sides?

ganeshie8 (ganeshie8):

Yes ! you got it !!

ganeshie8 (ganeshie8):

\(\large A = \frac{9}{2} (x + 23) \) multiply both sides wid 2 \(\large 2 \times A = 2 \times \frac{9}{2} (x + 23) \) \(\large 2A = 9(x + 23) \) divide 9 both sides \(\large \frac{2A}{9} = x + 23\) subtract 23 both sides \(\large \frac{2A}{9}-23 = x\)

OpenStudy (anonymous):

But then that would be what??? how do you subtract 23 from that fraction part thingy?? Thats so confusing!

ganeshie8 (ganeshie8):

you cant subtract, so leave it like that

ganeshie8 (ganeshie8):

we're done wid part A

OpenStudy (anonymous):

Oh wow... Thats confusing

ganeshie8 (ganeshie8):

:) get use to it

ganeshie8 (ganeshie8):

you wil see lot of these, u wil become comfortable wid these in no time... dont wry

OpenStudy (anonymous):

Okay lol I have no idea what the next question is even asking! >_<

ganeshie8 (ganeshie8):

put A = 108 above, and see wat u get

OpenStudy (anonymous):

so 2x108?

ganeshie8 (ganeshie8):

\(\large \frac{2A}{9}-23 = x\) put \(A = 108\) and find \(x\)

ganeshie8 (ganeshie8):

Yes ! \(\large \frac{2A}{9}-23 = x\) put \(A = 108\) \(\large \frac{2\times 108}{9}-23 = x\)

OpenStudy (anonymous):

okay so 216/9-23=x?

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

keep going

OpenStudy (anonymous):

24-23=x? so 1=x?

ganeshie8 (ganeshie8):

perfect ! 1 = x is same as x = 1

ganeshie8 (ganeshie8):

so the value of x when A = 108, is 1

ganeshie8 (ganeshie8):

can u do part C ?

OpenStudy (anonymous):

Okay thank you :) How about the last one?

ganeshie8 (ganeshie8):

you need to rip off things from n and isolate it

OpenStudy (anonymous):

No... How do you solve those?

OpenStudy (anonymous):

Okay...

ganeshie8 (ganeshie8):

give it a try... il help if u get stuck

ganeshie8 (ganeshie8):

-np - 90 > 30

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

so -np>120?

ganeshie8 (ganeshie8):

-90 is there on \(n\) side, so get rid of it first by 'adding 90 both sides'

ganeshie8 (ganeshie8):

thats right ! keep going

OpenStudy (anonymous):

I dunno what to do from there...

OpenStudy (anonymous):

add n to both sides>

ganeshie8 (ganeshie8):

No. you want \(n\)

OpenStudy (anonymous):

so add p?

ganeshie8 (ganeshie8):

since \(p\) is attached to \(n\), get rid of \(p\) 'by dividing \(p\) both sides'

ganeshie8 (ganeshie8):

-np>120 divide p both sides -n > 120/p

OpenStudy (anonymous):

is that my answer?

ganeshie8 (ganeshie8):

still n is not all by itself

ganeshie8 (ganeshie8):

you need to multiply -1 both sides to get rid of - sign

ganeshie8 (ganeshie8):

-n > 120/p n < -120/p

OpenStudy (anonymous):

How do I do that?

ganeshie8 (ganeshie8):

when u multiply NEGATIVE numbers, u need to flip signs

OpenStudy (anonymous):

oh so its just n<-120/p

ganeshie8 (ganeshie8):

yup

OpenStudy (anonymous):

Okay... Thank you!

ganeshie8 (ganeshie8):

yw !